# OVO-Bench: How Far is Your Video-LLMs from Real-World Online Video Understanding?

### v1

@@ Abstract
Temporal Awareness—the ability to reason dynamically based on the timestamp when a question is raised—is the key distinction between offline and online video LLMs. Unlike offline models, which rely on complete videos for static, post hoc analysis, online models process video streams incrementally and dynamically adapt their responses based on the timestamp at which the question is posed. Despite its significance, temporal awareness has not been adequately evaluated in existing benchmarks. To fill this gap, we present OVO-Bench (Online-VideO-Benchmark), a novel video benchmark that emphasizes the importance of timestamps for advanced online video understanding capability benchmarking. OVO-Bench evaluates the ability of video LLMs to reason and respond to events occurring at specific timestamps under three distinct scenarios: (1) Backward tracing: trace back to past events to answer the question. (2) Real-time understanding: understand and respond to events as they unfold at the current timestamp. (3) Forward active responding: delay the response until sufficient future information becomes available to answer the question accurately. OVO-Bench comprises 12 tasks, featuring 644 unique videos and approximately human-curated 2,800 fine-grained meta-annotations with precise timestamps. We combine automated generation pipelines with human curation. With these high-quality samples, we further developed an evaluation pipeline to systematically query video LLMs along the video timeline. Evaluations of nine Video-LLMs reveal that, despite advancements on traditional benchmarks, current models struggle with online video understanding, showing a significant gap compared to human agents. We hope OVO-Bench will drive progress in video LLMs and inspire future research in online video reasoning. Our benchmark and code can be accessed at https://github.com/JoeLeelyf/OVO-Bench.
@@ 1 Introduction
Building on the human-reviewed meta-annotations, we develop an evaluation pipeline that queries Video-LLMs densely along temporal axes to simulate continuous information processing. For Backward Tracing and Real-Time Visual Perception, we adopt multiple-choice evaluation, converting videos into segments from start to query time to accommodate offline models. With this approach, we explore the potential of explicitly leveraging state-of-the-art offline Video-LLMs for online video understanding. We evaluated 9 Video-LLMs, including proprietary models GPT-4o [35] and Gemini-1.5-Pro [45], alongside six recent open-source MLLMs like Qwen2-VL [47] and LLaVA-OneVision [22]. Despite their strong offline performance, these models struggle with online-style queries (e.g., What is happening now?), showing a significant gap from human performance. Further experiments on recent streaming models, such as Flash-VStream [58], reveal an even wider performance gap compared to offline counterparts, highlighting a substantial research space for further exploration and improvement.
@@ 2 Related Works
Online Video Understanding. Traditional offline video understanding methodologies primarily focus on accessing entire video sequences to facilitate prediction tasks. Conversely, online video understanding demands models to process video streams sequentially, making decisions based on current and past information. This approach is particularly well-suited for scenarios where future data is unavailable, such as in embodied intelligence, autonomous driving, and augmented reality applications. Among online video understanding methods, FlashVStream [58] employs a clustering method to select representative frames, enabling MLLMs for real-time interactions. LIVE [5] introduces a comprehensive framework for learning in video streams, which includes a training objective, data generation schema, and an inference pipeline tailored for online video understanding.
@@ 3.1.3 Forward Active Responding
Transitioning from passive reception to active perception is essential for advanced video understanding systems. Existing benchmarks primarily focus on the aforementioned two understanding modes, where Video-LLMs are required to respond immediately based on available information. In contrast, we introduce the Forward Active Responding mode, which allows the model to adjust its responses based on forthcoming visual inputs. We devise four task dimensions to evaluate the models’ active responding abilities:
@@ 3.2.1 Video and Annotation Collection
Meta-Annotations Collection. We employ three approaches to collect our meta-annotations which contain event-level timestamps: (1) Existing Annotation Repurposing. For human-annotated datasets with accurate event-level timestamps [1][44][15], we explicitly take advantage of these labels and reconstruct them to our final prompt. (3) Semi-Automatic Generation. For datasets that provide video-level QA pairs without complete temporal localization, including [32][50][38][20][13], we prompt temporal-sensitive Video-LLMs like Gemini-1.5[45] to provide coarse-grain timestamps which fit the event referred in question and answer. For tasks under the Real-Time Visual Perception scenario, timestamps are given during our automatic QA construction process, which will be illustrated in 3.2.2. We then perform meticulously inspect all collected source videos and the corresponding meta-annotations to ensure precision.
Figure 4: Left Queries Temporal Distribution in OVO-Bench. Center Linguistic Characteristics of Text Queries. Right Video category distribution of OVO-Bench.
@@ 3.3 Datasets Statistics
4 Experiments
4.1 Models and Evaluation Strategies
We evaluate four existing types of models: (1) Offline Multimodal Models, including GPT-4o [37], Gemini-1.5-Pro [45], Qwen2-VL [47], LLaVA-NeXT-Video [28], LLaVA-OneVision [27], InternVL-V2 [8] and LongVU [42], (2) Online Multimodal Models, including Flash-VStream-7B [58] and Videollm-Online[5] (3) Blind LLMs, including GPT-4-turbo [35]. (4) Human Agents. To ensure a fair comparison of model performance, we adhere to the principle of consistency by maintaining the same number of frames or frames per second (fps) across all models.
Considering the limitations on input video length for existing offline Video-LLMs, we adopt specialized video input methods tailored to such models. Specifically, we segment the video into clips based on the timestamps of the questions.For instance, for a question $Q_{i}$ posed at timestamp $t_{i}$, we extract the video clip $\text{Video}[0:t_{i}]$ as the visual input. This approach simulates a streaming question-answering scenario in online video understanding.
@@ 4.1 Models and Evaluation Strategies
Qwen2-VL-72B 64 72.48 56.88 77.59 52.25 74.26 61.41 65.81 51.52 73.65 63.44 62.87 37.68 60.10 45 47.59 58.76
LLaVA-NeXT-Video-7B 64 69.80 59.63 66.38 50.56 72.28 61.41 63.34 51.18 64.19 9.68 41.68 34.10 67.57 60.83 54.17 53.06
LLaVA-OneVision-7B 64 67.11 58.72 69.83 49.44 71.29 60.33 62.79 52.53 58.78 23.66 44.99 24.79 66.93 60.83 50.85 52.88
Qwen2-VL-7B 64 69.13 53.21 63.79 50.56 66.34 60.87 60.65 44.44 66.89 34.41 48.58 30.09 65.66 50.83 48.86 52.70
InternVL-V2-8B 64 68.46 58.72 68.97 44.94 67.33 55.98 60.73 43.10 61.49 27.41 44.00 25.79 57.55 52.92 45.42 50.05
LongVU-7B 1fps 55.70 49.54 59.48 48.31 68.32 63.04 57.40 43.10 66.22 9.14 39.49 16.62 69.00 60.00 48.54 48.48
@@
Flash-VStream-7B 1fps 25.50 32.11 29.31 33.71 29.70 28.80 29.86 36.36 33.78 5.91 25.35 5.44 67.25 60.00 44.23 33.15
VideoLLM-online-8B 2fps 8.05 23.85 12.07 14.04 45.54 21.20 20.79 22.22 18.80 12.18 17.73 - - - - -
Table 1: Detailed evaluation results on OVO-Bench. † refers to using ”low” resolution. To enhance the challenge of the questions by increasing the time interval between the question and the clues, the question time for [EPM] and [ASI] in the table is uniformly placed at the end of the video. [HLD] are evaluated in the standard manner following the streaming mode. For Forward Active Responding, accuracy-based evaluation metrics are utilized in this table. More high-quality [ASI] tasks are supplemented, further enhancing differentiation.
@@ 4.2 Main Results
Current Video-LLMs lack temporal prioritization when handling VQA tasks. Existing Video-LLMs do not prioritize real-time temporal information when answering questions, leading to an inability to accurately locate the correct scene when multiple misleading scenes matching the question appear in the video stream, as shown in Fig2. Even the best current proprietary models achieve only 54.49% and 66.97% on [STU] and [ACR] tasks, respectively, which represents a significant gap compared to Human Agents.
A powerful LLM backbone is the key to achieving high-performance video understanding. As shown in Table 1, the Qwen2-VL-72B model significantly outperforms its smaller counterpart, Qwen2-VL-7B, across all evaluated metrics. The larger model’s superior architecture allows it to excel in Real-Time Visual Perception with an average score of 65.81%, compared to 60.65% for the 7B variant. Additionally, Qwen2-VL-72B demonstrates enhanced capability in Backward Tracing tasks, achieving a score of 62.87%, notably higher than the 48.58% scored by Qwen2-VL-7B. These results highlight the importance of a robust and powerful LLM backbone in effectively processing and understanding complex video data, underscoring the need for more powerful architectures to achieve high performance in video understanding tasks.
@@ 6.1 Evaluation for Online Models on Forward Active Responding
Prompt Design. To adapt to the online scenarios, we constructed streaming mode prompts with accurate timestamps and also deleted the complicated instructional statement compared to 6.2. Prompts and examples of models’ responses are shown in 7.
@@ 9 Limitations
Offline Models for Online Video Understanding. As implied in our analysis 7, offline models usually perform random guesses in the forward active responding scenarios, making our evaluation unfair. For example, a model that always outputs ”Yes” can still achieve a score above zero in our evaluation. Moreover, the absence of online models with satisfactory performance, makes our benchmarks more suitable for future advancements. We hope our intensive work and intuitive ideas can guide the development of video understanding models toward real-world online video understanding.

### v2

@@ Abstract
Temporal Awareness—the ability to reason dynamically based on the timestamp when a question is raised—is the key distinction between offline and online video LLMs. Unlike offline models, which rely on complete videos for static, post hoc analysis, online models process video streams incrementally and dynamically adapt their responses based on the timestamp at which the question is posed. Despite its significance, temporal awareness has not been adequately evaluated in existing benchmarks. To fill this gap, we present OVO-Bench (Online-VideO-Benchmark), a novel video benchmark that emphasizes the importance of timestamps for advanced online video understanding capability benchmarking. OVO-Bench evaluates the ability of video LLMs to reason and respond to events occurring at specific timestamps under three distinct scenarios: (1) Backward tracing: trace back to past events to answer the question. (2) Real-time understanding: understand and respond to events as they unfold at the current timestamp. (3) Forward active responding: delay the response until sufficient future information becomes available to answer the question accurately. OVO-Bench comprises 12 tasks, featuring 644 unique videos and approximately human-curated 2,800 fine-grained meta-annotations with precise timestamps. We combine automated generation pipelines with human curation. With these high-quality samples, we further developed an evaluation pipeline to systematically query video LLMs along the video timeline. Evaluations of eleven Video-LLMs reveal that, despite advancements on traditional benchmarks, current models struggle with online video understanding, showing a significant gap compared to human agents. We hope OVO-Bench will drive progress in video LLMs and inspire future research in online video reasoning. Our benchmark and code can be accessed at https://github.com/JoeLeelyf/OVO-Bench.
@@ 1 Introduction
Building on the human-reviewed meta-annotations, we develop an evaluation pipeline that queries Video-LLMs densely along temporal axes to simulate continuous information processing. For Backward Tracing and Real-Time Visual Perception, we adopt multiple-choice evaluation, converting videos into segments from start to query time to accommodate offline models. With this approach, we explore the potential of explicitly leveraging state-of-the-art offline Video-LLMs for online video understanding. We evaluated eleven Video-LLMs, including proprietary models GPT-4o [34] and Gemini-1.5-Pro [44], alongside six recent open-source MLLMs like Qwen2-VL [46] and LLaVA-OneVision [22]. Despite their strong offline performance, these models struggle with online-style queries (e.g., What is happening now?), showing a significant gap from human performance. Further experiments on recent streaming models, such as Flash-VStream [57], reveal an even wider performance gap compared to offline counterparts, highlighting a substantial research space for further exploration and improvement.
@@ 2 Related Works
Online Video Understanding. Traditional offline video understanding methodologies primarily focus on accessing entire video sequences to facilitate prediction tasks. Conversely, online video understanding demands models to process video streams sequentially, making decisions based on current and past information. This approach is particularly well-suited for scenarios where future data is unavailable, such as in embodied intelligence, autonomous driving, and augmented reality applications. Among online video understanding methods [60, 38], FlashVStream [57] employs a clustering method to select representative frames, enabling MLLMs for real-time interactions. LIVE [5] introduces a comprehensive framework for learning in video streams, which includes a training objective, data generation schema, and an inference pipeline for online video understanding.
@@ 3.1.3 Forward Active Responding
Transitioning from passive reception to active perception is essential for advanced video understanding systems. Existing benchmarks primarily focus on the aforementioned two understanding modes, where Video-LLMs are required to respond immediately based on available information. In contrast, we introduce the Forward Active Responding mode, which allows the model to adjust its responses based on forthcoming visual inputs. We devise three task dimensions to evaluate the models’ active responding abilities:
@@ 3.2.1 Video and Annotation Collection
Meta-Annotations Collection. We employ three approaches to collect our meta-annotations, which contain event-level timestamps: (1) Existing Annotation Repurposing. For human-annotated datasets with accurate event-level timestamps [1][43][15], we explicitly take advantage of these labels and reconstruct them to our final prompt. (2) Semi-Automatic Generation. For datasets that provide video-level QA pairs without complete temporal localization, including [31][49][36][20][13], we prompt temporal-sensitive Video-LLMs like Gemini-1.5[44] to provide coarse-grain timestamps which fit the event referred in question and answer. For tasks under the Real-Time Visual Perception scenario, timestamps are given during our automatic QA construction process, which will be illustrated in 3.2.2. (3) Human-annotated. For the [SSR] and [CRR] tasks, questions, answers, and ground-truth timestamps are collected by our recruited volunteers. We then perform a meticulous inspection of all collected source videos and the corresponding meta-annotations to ensure precision.
Figure 4: Left: Queries Temporal Distribution in OVO-Bench. Center: Linguistic Characteristics of Text Queries. Right: Video category distribution of OVO-Bench.
@@ 3.3 Datasets Statistics
Qwen2-VL-72B[46] 64 65.77 60.55 69.83 51.69 69.31 54.35 61.92 52.53 60.81 57.53 56.95 38.83 64.07 45.00 49.30 56.27
LLaVA-Video-7B[61] 64 69.13 58.72 68.83 49.44 74.26 59.78 63.52 56.23 57.43 7.53 40.4 34.10 69.95 60.42 54.82 52.91
LLaVA-OneVision-7B[22] 64 66.44 57.80 73.28 53.37 71.29 61.96 64.02 54.21 55.41 21.51 43.71 25.64 67.09 58.75 50.50 52.74
Qwen2-VL-7B[46] 64 60.40 50.46 56.03 47.19 66.34 55.43 55.98 47.81 35.48 56.08 46.46 31.66 65.82 48.75 48.74 50.39
InternVL-V2-8B[8] 64 67.11 60.55 63.79 46.07 68.32 56.52 60.39 48.15 57.43 24.73 43.44 26.5 59.14 54.14 46.60 50.15
LongVU-7B[41] 1fps 53.69 53.21 62.93 47.75 68.32 59.78 57.61 40.74 59.46 4.84 35.01 12.18 69.48 60.83 47.50 46.71
@@
Flash-VStream-7B[57] 1fps 24.16 29.36 28.45 33.71 25.74 28.80 28.37 39.06 37.16 5.91 27.38 8.02 67.25 60.00 45.09 33.61
VideoLLM-online-8B[5] 2fps 8.05 23.85 12.07 14.04 45.54 21.20 20.79 22.22 18.80 12.18 17.73 - - - - -
Dispider[38] 1fps 57.72 49.54 62.07 44.94 61.39 51.63 54.55 48.48 55.41 4.3 36.06 18.05 37.36 48.75 34.72 41.78
Table 1: Detailed evaluation results on OVO-Bench. To enhance the challenge of the questions by increasing the time interval between the question and the clues, the question time for [EPM] and [ASI] in the table is uniformly placed at the end of the video. For Forward Active Responding, accuracy-based evaluation metrics are utilized in this table.
4 Experiments
4.1 Models and Evaluation Strategies
We evaluate four existing types of models: (1) Offline Multimodal Models, including GPT-4o [35], Gemini-1.5-Pro [44], Qwen2-VL [46], LLaVA-Video [61], LLaVA-OneVision [27], InternVL-V2 [8] and LongVU [41], (2) Online Multimodal Models, including FlashVStream [57] , Videollm-Online[5] and Dispider[38], (3) Blind LLMs, including GPT-4-turbo [34]. (4) Human Agents. To ensure a fair comparison of model performance, we adhere to the principle of consistency by maintaining the same number of frames or frames per second (fps) across all models.
Considering the input video length limitations for offline Video-LLMs, we adopt specialized video input methods tailored to such models. Specifically, we segment the video into clips based on the timestamps of the questions.For instance, for a question $Q_{i}$ posed at timestamp $t_{i}$, we extract the video clip $\text{Video}[0:t_{i}]$ as the visual input. This approach simulates a streaming question-answering scenario in online video understanding.
We also conduct a runtimes study of five models, including QWen2-VL-7B [46], LLaVA-Video [61], LLaVA-OneVision [22], InternVL-V2[8], and FlashVStream [57]. In this setting, we randomly select 100 samples from tasks in Backward Tracing and Real-Time Visual Perception and then plot the change of average inference delay on these videos with the number of sampled frames.
@@ 4.2 Main Results
Current Video-LLMs lack temporal prioritization when handling VQA tasks. Existing Video-LLMs do not prioritize real-time temporal information when answering questions, leading to an inability to accurately locate the correct scene when multiple misleading scenes matching the question appear in the video stream, as shown in Fig2. Even the best current proprietary models achieve only 58.43% and 66.97% on [STU] and [ACR] tasks, respectively, which represents a significant gap compared to Human Agents.
@@
Current Video-LLMs need more efficient inference frameworks to achieve real-time visual question answering. As shown in Fig6, the inference latencies of current Video-LLMs exhibit an exponential growth trend as frame numbers increase. Specifically, when using 64 frames as visual input, most efficient Video-LLMs, like QWen2VL-7B[46] and FlashVStream [57], still need around 4 seconds on average to perform a response, making real-time video dialogue far from reach.
@@ 4.4 Forward Active Responding
Figure 6: Inference Latency (y-axis) v.s. Frames Number (x-axis). Latency test on an A100 GPU for FlashVStream and four A100 GPUs for other models.
@@ 5 Conclusion and Future Work
Acknowledgement
This project is funded in part by Shanghai Artificial lntelligence Laboratory, Shanghai Innovation Institute, the National Key R&D Program of China (2022ZD0160201). This project is also supported by the Shanghai Postdoctoral Excellence Program (No.2023023), China Postdoctoral Science Fund (No.2024M751559).
[…]
@@ 6.1 Evaluation for Online Models on Forward Active Responding
Prompt Design. To adapt to the online scenarios, we constructed streaming mode prompts with accurate timestamps and also deleted the complicated instructional statement compared to 6.2. Prompts and examples of models’ responses are shown in 8.
@@ 9 Limitations
Offline Models for Online Video Understanding. As implied in our analysis 8, offline models usually perform random guesses in the forward active responding scenarios, making our evaluation unfair. For example, a model that always outputs ”Yes” can still achieve a score above zero in our evaluation. Moreover, the absence of online models with satisfactory performance, makes our benchmarks more suitable for future advancements. We hope our intensive work and intuitive ideas can guide the development of video understanding models toward real-world online video understanding.